OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 8 and a second term of 5?

OpenStudy (anonymous):

is it multiple choice

OpenStudy (tkhunny):

We need a1 - The first term We need d - The common difference.

OpenStudy (anonymous):

@gabbimanges17

OpenStudy (anonymous):

an = 8 − 5(n − 1); all integers where n ≥ 1 an = 8 − 5(n − 1); all integers where n ≥ 0 an = 8 − 3(n − 1); all integers where n ≥ 1 an = 8 − 3(n − 1); all integers where n ≥ 0

OpenStudy (anonymous):

@_greatmath7

OpenStudy (tkhunny):

How much did the value change from the first value to the second value?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

First term = 8 Common difference = 5−8 = −3 an = 8 − 3(n − 1); all integers where n ≥ 1

OpenStudy (tkhunny):

Good. Throw out the two that say "-5".

OpenStudy (anonymous):

-5

OpenStudy (tkhunny):

That's right. Throw those out. The common difference is -3. A formula with "-5" can't do that. Scratch them out and never look at them again. Focus on the ones with "-3".